OpenStudy (anonymous):

summation of n root n/(n+1)^2 form 1 to infinity

OpenStudy (anonymous):

is that a question or an affirmation?

OpenStudy (anonymous):

QUESTION. DOES IT DIVERGE OR CONVERGE AND WHAT TEST do i use'

OpenStudy (anonymous):

this does not converge for sure

OpenStudy (anonymous):

Its a question, i'm not sure where to start. sorry my keyboard is messed up

OpenStudy (anonymous):

what part of math is it? I think you can plut numbers from 0 to 10 (for example) and see how it converges

OpenStudy (anonymous):

if it is \[\sum_1^{\infty}\frac{n\sqrt{n}}{(n+1)^2}\]

OpenStudy (anonymous):

is that the question?

OpenStudy (anonymous):

its calculus 2 infinite series and thats it!!

OpenStudy (anonymous):

does not converge. we do it with our eyeballs.

OpenStudy (anonymous):

the denominator is a polynomial of degree 2. the numerator has "degree" 3/2

OpenStudy (anonymous):

in order for this to converge the degree of the denominator has to be greater than one more than the degree of the numerator

OpenStudy (anonymous):

so you subtract the numerator degree and denominator degree and use the p test to determine convergence?

OpenStudy (anonymous):

yes that would work

OpenStudy (anonymous):

the degree of the numerator is 1/n tho

OpenStudy (anonymous):

also probably comparison test because this is evidently bigger than \[\sum\frac{1}{n}\] which diverges

OpenStudy (anonymous):

?? \[n\times \sqrt{n}=n^1\times n^{\frac{1}{2}}=n^{\frac{3}{2}}\]

OpenStudy (anonymous):

wait since n^1/n-2<0 for all n...that would make p<1. wouldn't that diverge?

OpenStudy (anonymous):

just to help prove your point...I just wanna know if that's is that a correct proof

OpenStudy (anonymous):

i am afraid you have lost me. the denominator is a polynomial of degree 2. it is \[n^2+2n+1\] the numerator is \[n^{\frac{3}{2}}\] so even without the \[\sqrt{n}\] part this would diverge. you have \[\frac{n^{\frac{3}{2}}}{n^2+2n+1}\]

OpenStudy (anonymous):

so with my eyeballs i say \[2-\frac{3}{2}=\frac{1}{2}\] and \[\frac{1}{2}<1\] so by comparison test (to say 1/n) this diverges

OpenStudy (anonymous):

maybe you should write something fancier, but i am not sure what.

OpenStudy (anonymous):

i see u must have multiplied both the numerator and denominator by n. I just theorized \[\sum_{?}^{?} n^1/n \div n^2 = \sum_{?}^{?} n^1/n-2\] plus any number form 1 to infinity in for n and the number will always be less than one. Thanks for your help!!

OpenStudy (anonymous):

plug*

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

I'm wrong?